OpenStudy (anonymous):

The equation of line LM is y = 5x + 4. Write an equation of a line perpendicular to line LM in slope-intercept form that contains point (-3, 2).

OpenStudy (anonymous):

Select one: a. y = -5x + 13 b. y=-1/5x + 7/5 c. y=1/5x - 7/5 d. y=5x- 17

OpenStudy (anonymous):

@Nurali

OpenStudy (luigi0210):

Perpendicular means the negative reciprocal.

OpenStudy (anonymous):

ok

OpenStudy (luigi0210):

So the slope becomes: \[\LARGE 5 ===> -\frac{1}{5}\]

OpenStudy (luigi0210):

And \[\LARGE y=mx+b\] m=slope b=x-intercept

OpenStudy (anonymous):

isn't that the slope formula

OpenStudy (luigi0210):

Yup, so if you plug in -1/5 you have.. \[\LARGE y=-\frac{1}{5}x+b\] And there's only one answer choice with that slope..

OpenStudy (anonymous):

so its b or c

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(-3\quad ,&2) \end{array} \\\quad \\ slope = m= -\cfrac{1}{5} \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)

OpenStudy (luigi0210):

Made a little mistake, the b is the y-intercept, not x

OpenStudy (luigi0210):

Well, c has a POSITIVE slope, we need the one with the negative slope